A & B are undergoing first order rxn with rate constant K1 & K2 which are in the ratio 3:2. If initial concentration of A & B are in ratio 3:2, what would be ratio of [A] & [B] after three half lifes of A ?

Grade:12th pass

2 Answers

Gopesh Kumar IIT Roorkee
9 years ago
k_1=1.5\times k_2
So, T_{1/2}\left ( 1 \right )=(2/3)T_{1/2}\left ( 2 \right ) Therefore, 3T_{1/2}\left ( 1 \right )=2T_{1/2}\left ( 2 \right ) Hence, [B]=0.25[B]_{0} and [A]=0.125[A]_{0} Calculating the ratio, we get, the answer as 3:4.
Arpit Dhankar
9 years ago
thank u sir!